Jane leaves her house and drives in a straight line at a speed of 39 mph. Her speed on the return trip is 49 mph. The trip back home takes 15 minutes less than the outbound trip. What was Jane's total travel time in hours? Assume both trips are in a straight line. Give your answer to 2 decimal places
Answer by richwmiller(17219)

r*t=d
the distance is the same in both directions.
39*t=49(t-1/4)
39t=49t-49/4
0=10t-49/4
49/4=10t
4.9/4=t
1.225=t
what is the rule for rounding ?
Is it 1.22 or 1.23?